# The Coding Limits of Robust Watermarking for Generative Models

### v1

@@ 2.2 Watermarking for Generative Models
We recall the formalization of watermarking for generative models, following [ZEF+23, ZEF+24].
@@
A conditional generative model $\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}$ is a probabilistic polynomial-time (PPT) algorithm that, given a prompt $x\in\mathcal{X}$, outputs $y\in\mathcal{Y}$. Here, $\mathcal{X}$ is the prompt space, and $\mathcal{Y}$ is the output space. We write $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Model}(x)$ to denote sampling a response from the model on prompt $x$.
@@
A robust watermarking scheme should satisfy three key properties. First, correctness: watermarked outputs should always be recognized as such. Second, soundness: non-watermarked outputs (such as human-written or independently generated model outputs) should almost never be mistakenly detected as watermarked. Third, robustness: the watermark should survive small modifications to the output. That is, an adversary cannot erase the watermark by making limited edits unless a large fraction of the output is changed.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}\}$ be a class of generative models. We say that $\Pi$ satisfies correctness if for every $\mathsf{Model}\in\mathcal{M}$ and for every $x\in\mathcal{X}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,y)=\mathtt{true}:(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Watermark}(1^{\lambda},\mathsf{Model}),\ y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Model}_{\kappa}(x)}}\right]=1$.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}\}$ be a class of generative models. We say that $\Pi$ satisfies soundness if for every $\mathsf{Model}\in\mathcal{M}$, every $x\in\mathcal{X}$, and every $y\in\mathcal{Y}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,y)=\mathtt{true}:(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Watermark}(1^{\lambda},\mathsf{Model})}}\right]\leq\mathsf{negl}(\lambda)$.
Robustness requires that a watermark remain detectable even if an adversary modifies a small part of the output. To capture this, we consider tampering functions $f:\Sigma^{n}\rightarrow\Sigma^{n}$ which, for a fixed parameter $\alpha\in(0,1)$, output a string of the same length as the input and differ from the original output $y\in\Sigma^{n}$ in at most $\alpha n$ positions. Here, $\alpha$ specifies the maximum allowable fraction of edits.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\Sigma^{*}\rightarrow\Sigma^{*}\}$ be a class of generative models. A watermarking scheme $\Pi$ is $^\hat{\mathcal{F}}$-robust if for every $\mathsf{Model}\in\mathcal{M}$, every $x\in\mathcal{X}$, and every $f\in\hat{\mathcal{F}}_{\alpha}$,
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,\tilde{y})=\mathtt{false}\land\tilde{y}\neq y:\begin{array}[]{c}(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Watermark}(1^{\lambda},\mathsf{Model})\\ y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Model}_{\kappa}(x);\;\tilde{y}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}f(y)\end{array}}}\right]\leq\mathsf{negl}(\lambda).$
In our results, we primarily consider small constants $\alpha>0$. When we refer to $\hat{\mathcal{F}}^{\mathrm{ind}}_{\alpha}$, the impossibility results become even stronger: if robustness fails against independent (random, uncoordinated) edits, it necessarily fails against arbitrary (coordinated) ones.
@@
3 Secret-key Coding Schemes
In this section, we develop a central ingredient for our theory: secret-key codes that do not encode information, but instead provide a way to test the validity of a codeword and to detect tampering. These “messageless” codes distill the core cryptographic challenges of watermarking into a simple form. We formalize their properties, such as correctness, soundness, tamper detection, and pseudorandomness, and see how these definitions serve as the foundation for our results on watermarking.
3.1 Messageless Secret-key Codes
We begin by defining messageless secret-key codes: codes whose main functionality is to enable (i) distinguishing valid codewords from invalid ones, and (ii) detecting whether a codeword has been tampered with (possibly maliciously) during transmission. We consider codes in the secret-key setting, where both encoding and decoding use the same secret key. This secret-key variant is also simpler to construct than public-key analogues.
Formally, a messageless secret-key code is specified by a triple of polynomial-time algorithms $\Gamma=(\mathsf{KGen},\mathsf{Enc},\mathsf{Dec})$ over an alphabet $\Sigma$} and a codeword space ${\Sigma}^{n}$:
@@ 3.1 Messageless Secret-key Codes
On input the security parameter $1^{\lambda}$}, the randomized key-generation algorithm outputs a secret key $\mathit{sk}$.
@@
On input $\mathit{sk}$, the (possibly randomized) encoding algorithm outputs a codeword $\gamma$.
@@
On input the secret key $\mathit{sk}$ and a codeword $\gamma$, the deterministic decoding algorithm outputs a symbol $y\in\{\mathtt{valid},\mathtt{invalid},\mathtt{tampered}\}$.
A messageless secret-key code $\Gamma$ should satisfy correctness: an honestly generated codeword is always classified as $\mathtt{valid}$ by the decoding procedure.
Definition 5 (Correctness of messageless secret-key codes).
We say that a messageless secret-key code $\Gamma$ satisfies correctness if for all $\lambda\in\mathbb{N}$ and for all $\mathit{sk}\in\mathsf{KGen}(1^{\lambda})$, it holds that $\mathsf{Dec}(\mathit{sk},\mathsf{Enc}(\mathit{sk}))=\mathtt{valid}$. If $\mathsf{Enc}$ is randomized, we require $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\mathsf{Enc}(\mathit{sk}))=\mathtt{valid}}}\right]=1$ where the probability is over the randomness of the encoding.
For security, we require two properties. The first is soundness, which means that decoding any fixed string (not produced by the encoder) yields $\mathtt{invalid}$ except with negligible probability over the choice of the secret key $\mathit{sk}$. In other words, the code identifies codewords associated to $\mathit{sk}$.
Definition 6 (Soundness of messageless secret-key codes).
@@
We say that a messageless secret-key code $\Gamma$ satisfies soundness if for every fixed codeword $\hat{\gamma}\in{\Sigma}^{n}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]\leq\mathsf{negl}(\lambda).$
The second property is tamper detection with respect to a family $\mathcal{F}=\{f:{\Sigma}^{n}\rightarrow{\Sigma}^{n}\}$ of codeword modification functions (possibly randomized). This property requires that decoding a tampered codeword $\tilde{\gamma}\neq\gamma$ (with $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Enc}(\mathit{sk})$ and $\tilde{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}f(\gamma)$) yields $\mathtt{tampered}$ except with negligible probability.
Definition 7 ($\mathcal{F}$-Tamper detection of messageless secret-key codes).
We say a messageless secret-key code $\Gamma$ satisfies $\mathcal{F}$-tamper detection if for any $f\in\mathcal{F}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mathtt{tampered}\wedge\tilde{\gamma}\neq\gamma:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda});\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Enc}(\mathit{sk});\tilde{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}f(\gamma)}}\right]\leq\mathsf{negl}(\lambda).$
Finally, in some settings, we require a third property: pseudorandomness. Intuitively, a messageless secret-key code is pseudorandom if the output of $\mathsf{Enc}$ is indistinguishable from a uniformly random string of the same length, even if $\mathsf{Enc}$ is called multiple times. This property requires that $\mathsf{Enc}$ is randomized (since a deterministic encoder is trivially distinguishable from random).
Definition 8 (Pseudorandomness of messageless secret-key codes).
We say that a messageless secret-key code $\Gamma$ satisfies pseudorandomness if for every PPT distinguisher $\mathsf{D}$,
@@
$\mathopen{}\mathclose{{\left\lvert\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{D}^{\mathcal{O}_{\sf real}}(1^{\lambda})=1:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]-\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{D}^{\mathcal{O}_{\sf rand}}(1^{\lambda})=1}}\right]}}\right\rvert\leq\mathsf{negl}(\lambda),$
where the inputless oracle $\mathcal{O}_{\sf real}$ (resp. $\mathcal{O}_{\sf rand}$), each time it is invoked, returns $\mathsf{Enc}(\mathit{sk})$ (resp. $\rho{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}{\Sigma}^{n}$).

### v2

@@ 2.2 Watermarking for Generative Models
We recall the formalization of watermarking for generative models, following [zhang2023watermarks, zhang2024watermarks].
@@
A conditional generative model $\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}$ is a probabilistic polynomial-time (PPT) algorithm that, given a prompt $x\in\mathcal{X}$, outputs $y\in\mathcal{Y}$. Here, $\mathcal{X}$ is the prompt space, and $\mathcal{Y}$ is the output space. We write $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Model}(x)$ to denote sampling a response from the model on prompt $x$.
@@
A robust watermarking scheme should satisfy three key properties. First, correctness: watermarked outputs should always be recognized as such.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}\}$ be a class of generative models. We say that $\Pi$ satisfies correctness if for every $\mathsf{Model}\in\mathcal{M}$ and for every $x\in\mathcal{X}$, we have
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,y)=\mathtt{true}:(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Watermark}(1^{\lambda},\allowbreak\mathsf{Model}),\ y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Model}_{\kappa}(x)}}\right]=1.$
Second, soundness: non-watermarked outputs (such as human-written or independently generated model outputs) should almost never be mistakenly detected as watermarked.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\mathcal{X}\rightarrow\mathcal{Y}\}$ be a class of generative models. We say that $\Pi$ satisfies soundness if for every $\mathsf{Model}\in\mathcal{M}$, every $x\in\mathcal{X}$, and every $y\in\mathcal{Y}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,y)=\mathtt{true}:(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Watermark}(1^{\lambda},\mathsf{Model})}}\right]\leq\mathsf{negl}(\lambda)$.
Third, robustness: the watermark should survive small modifications to the output. Robustness requires that a watermark remain detectable even if an adversary modifies a small part of the output. To capture this, we consider tampering functions $f:\Sigma^{n}\rightarrow\Sigma^{n}$ which, for a fixed parameter $\alpha\in(0,1)$, output a string of the same length as the input and differ from the original output $y\in\Sigma^{n}$ in at most $\alpha n$ positions. Here, $\alpha$ specifies the maximum allowable fraction of edits.
@@
Let $\mathcal{M}=\{\mathsf{Model}:\Sigma^{*}\rightarrow\Sigma^{*}\}$ be a class of generative models. A watermarking scheme $\Pi$ is $^\hat{\mathcal{F}}$-robust if for every $\mathsf{Model}\in\mathcal{M}$, every $x\in\mathcal{X}$, and every $f\in\hat{\mathcal{F}}_{\alpha}$, we have
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Detect}(\kappa,x,\tilde{y})=\mathtt{false}\land\tilde{y}\neq y:\begin{array}[]{c}(\kappa,\mathsf{Model}_{\kappa}){\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Watermark}(1^{\lambda},\mathsf{Model})\\ y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Model}_{\kappa}(x);\;\tilde{y}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}f(y)\end{array}}}\right]\leq\mathsf{negl}(\lambda).$
In our results, we primarily consider small constants $\alpha>0$. When we refer to $\hat{\mathcal{F}}^{\mathrm{ind}}_{\alpha}$, the impossibility results are strongest, in the sense that we show robustness is already impossible against this restricted class of independent, uncoordinated edits. This immediately implies impossibility against the larger class of arbitrary tampering functions.
@@
3 Zero-bit Tamper-detection Codes
Our lower bounds will be phrased in terms of a simple coding primitive that captures the behavior of watermark detectors. Informally, we imagine that for each secret key there is a single special codeword in ${\Sigma}^{n}$. Given a string $y\in{\Sigma}^{n}$, the decoder must decide whether
  1. 1.
    $y$ is the honest codeword produced under the key,
  2. 2.
    $y$ is unrelated to any codeword for that key, or
  3. 3.
    $y$ should be treated as a tampered version of the honest codeword.
There is no explicit message to recover; the output is only this three-way classification. We refer to such objects as zero-bit tamper-detection codes.
We work in a secret-key setting: both the encoder and the decoder have access to the same key $\mathit{sk}$, which is hidden from the adversary. This is a favorable model for the designer of the code. Any public-key or keyless detector can be viewed as a special case in which the key is fixed and revealed in advance. Thus, if we show that no scheme exists even in this more permissive secret-key model, the impossibility automatically applies to all more restricted variants.
It is helpful to compare this notion with a familiar primitive. Suppose we have a secure message authentication code with tagging algorithm $\mathsf{Tag}_{\mathit{sk}}(\cdot)$. Fix a message $m=0$, and for a given key $\mathit{sk}$ let $\gamma=\mathsf{Tag}_{\mathit{sk}}(0)$. We can define an encoder that always outputs $\gamma$, and a decoder that declares $\mathtt{valid}$ on input $\gamma$ and $\mathtt{tampered}$ on every other input. This construction detects any change to $\gamma$ perfectly. In this sense, detecting modifications of a single designated string is easy. The difficulty arises once we also require the decoder to behave well on all other strings that might be presented to it. The definitions below make this requirement precise.
Formally, a zero-bit tamper-detection code is specified by a triple of polynomial-time algorithms
$\Gamma=(\mathsf{KGen},\mathsf{Enc},\mathsf{Dec})$
over an alphabet $\Sigma$} and a codeword space ${\Sigma}^{n}$:
@@ 3 Zero-bit Tamper-detection Codes
On input the security parameter $\lambda$, the randomized key-generation algorithm outputs a secret key $\mathit{sk}$.
@@
On input $\mathit{sk}$, the (possibly randomized) encoding algorithm outputs a codeword $\gamma\in{\Sigma}^{n}$.
@@
On input the secret key $\mathit{sk}$ and a string $\gamma\in{\Sigma}^{n}$, the deterministic decoding algorithm outputs a symbol in $\{\mathtt{valid},\mathtt{invalid},\mathtt{tampered}\}$.
The first requirement is that honestly generated codewords are always accepted.
Definition 5 (Correctness of zero-bit tamper-detection codes).
A zero-bit tamper-detection code $\Gamma$ satisfies correctness if, for every $\lambda\in\mathbb{N}$ and every secret key $\mathit{sk}\in\mathrm{supp}(\mathsf{KGen}(1^{\lambda}))$, we have
$\mathbb{P}\bigl[\mathsf{Dec}(\mathit{sk},\mathsf{Enc}(\mathit{sk}))=\mathtt{valid}\bigr]=1,$
where the probability is taken over the internal randomness of $\mathsf{Enc}$ (if any). In particular, if $\mathsf{Enc}$ is deterministic then $\mathsf{Dec}(\mathit{sk},\mathsf{Enc}(\mathit{sk}))=\mathtt{valid}$ holds for all such $\mathit{sk}$.
To capture security, we add two further conditions. The first is soundness. Here we fix an arbitrary string $\hat{\gamma}\in{\Sigma}^{n}$ that is chosen independently of the key. Soundness requires that, over the random choice of $\mathit{sk}$, the decoder almost never classifies $\hat{\gamma}$ as anything other than $\mathtt{invalid}$. Intuitively, for a random key the set of strings that the decoder considers valid or tampered should occupy only a small fraction of the ambient space.
Definition 6 (Soundness of zero-bit tamper-detection codes).
@@
We say that a zero-bit tamper-detection code $\Gamma$ satisfies soundness if for every fixed string $\hat{\gamma}\in{\Sigma}^{n}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]\leq\mathsf{negl}(\lambda).$
For our lower bounds we will also use a relaxed form of this requirement. Rather than insisting that the misclassification probability be negligible, we allow it to be bounded by a fixed constant $c$. From the point of view of applications, such a guarantee is very weak. For example, if $c=1/2$, then in the worst case the decoder can do no better than a fair coin flip on a fixed string. The relaxation is useful for the structure of our result: the class of schemes that satisfy negligible soundness is contained inside the larger class that satisfy constant-bounded soundness for some $c<1$. If we can rule out robust tamper detection for the larger class, then we have automatically ruled it out for the smaller class as well.
Definition 7 (Constant-bounded soundness of zero-bit tamper-detection codes).
Let $c\in(0,1]$. We say that a zero-bit tamper-detection code $\Gamma$ satisfies constant-bounded soundness (with parameter $c$) if for every fixed string $\hat{\gamma}\in{\Sigma}^{n}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]\leq c.$
If $c=\tfrac{1}{2}$, we say that $\Gamma$ satisfies coin-flipping soundness, in the sense that checking the validity of a worst-case string is no better than flipping a fair coin.
The second security condition concerns the effect of tampering on an honestly generated codeword. We model tampering by a family $\mathcal{F}=\{f:{\Sigma}^{n}\rightarrow{\Sigma}^{n}\}$ of (possibly randomized) functions. For a given key, the encoder produces $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Enc}(\mathit{sk})$ and the adversary applies some $f\in\mathcal{F}$ to obtain $\tilde{\gamma}$. Tamper detection requires that, except with negligible probability, the decoder outputs $\mathtt{tampered}$ whenever $\tilde{\gamma}$ differs from $\gamma$.
Definition 8 ($\mathcal{F}$-tamper detection of zero-bit tamper-detection codes).
We say a zero-bit tamper-detection code $\Gamma$ satisfies $\mathcal{F}$-tamper detection if for every $f\in\mathcal{F}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mathtt{tampered}\wedge\tilde{\gamma}\neq\gamma:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda});\;\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Enc}(\mathit{sk});\;\tilde{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}f(\gamma)}}\right]\leq\mathsf{negl}(\lambda).$
Correctness, soundness, and $\mathcal{F}$-tamper detection together describe the behavior we will require from the coding layer underlying cryptographic watermarking.
3.1 Relation to PRCs
Christ and Gunn [C:ChrGun24] formalize cryptographic watermarking in terms of pseudorandom error-correcting codes (PRCs). A PRC is a keyed coding scheme that both corrects a bounded amount of noise and hides its codewords from anyone who does not know the key. Since our lower bounds are meant to apply to PRC-based watermarking, it is natural to ask why we did not state them directly in that language.
The main reason is that we work in the information theoretic setting to prove the negative result. [^1] Despite this, the negative result can be naturally extended even to computational security. It depends only on how many symbols an adversary can change and on how the decoder responds to corrupted strings. The pseudorandomness of the codeword, which is central to undetectability, plays no role in this counting argument. Zero-bit tamper-detection codes isolate exactly the pieces we need: a keyed encoder, a decoder that distinguishes $\mathtt{valid}$, $\mathtt{invalid}$, and $\mathtt{tampered}$, and quantitative guarantees for soundness and tamper detection with respect to a family $\mathcal{F}$ of tampering functions. We do not constrain the distribution of codewords beyond what is needed for these properties.
Viewed this way, zero-bit tamper-detection codes form a weaker abstraction than PRCs. A zero-bit PRC must in addition make its codewords pseudorandom and is usually required to have negligible decoding error for every message. Any impossibility theorem that already rules out robust zero-bit tamper detection under these minimal assumptions therefore applies a fortiori to zero-bit PRCs that satisfy stronger guarantees. Working with the weaker interface keeps the statement of the lower bound focused on the two properties that matter for our application (soundness and tamper detection) while still covering the PRC constructions used in cryptographic watermarking.
At the same time, in the zero-bit setting the coding content of the two formalisms is essentially the same once we impose standard soundness. There is only a single message $\mu$ to recover. Given a zero-bit tamper-detection code $\Gamma=(\mathsf{KGen},\mathsf{Enc},\mathsf{Dec})$ that satisfies soundness in the sense of 6 and $\mathcal{F}$-tamper detection in the sense of 8, we can view it as a (non-pseudorandom) zero-bit error-correcting code for the fixed message $\mu$ by using $\mathsf{Enc}$ as the encoder and defining the decoder to output $\mu$ on any input that is not classified as $\mathtt{invalid}$. Conversely, any zero-bit PRC with negligible decoding error in a given tampering model induces a tamper-detection code by interpreting successful decodings as $\mathtt{valid}$ and all other outcomes as $\mathtt{tampered}$. In both directions, the soundness and tamper-detection parameters are preserved up to the usual negligible losses.
Thus, if we strip away pseudorandomness, zero-bit tamper-detection codes and zero-bit PRCs express the same coding-theoretic limitations in the regime we study. Our lower bound for zero-bit tamper-detection codes immediately carries over to zero-bit PRCs whose decoder is required to be robust against the same independent symbol tampering model. Adding pseudorandomness on top can improve undetectability of the watermark, but it cannot move the underlying coding limit.
Finally, the limit we obtain is tight for this abstraction. In Appendix A we describe a simple information-theoretic construction of a zero-bit tamper-detection code that tolerates any independent tampering rate strictly below the threshold identified by our bound. This matching construction shows that the impossibility is not an artifact of the definition: it captures the exact corruption level at which robust decoding, and hence robust PRC-based watermarking, becomes impossible.